# Chirality induced phase separation in active circle swimmers

### v1

@@ III.1 Steady state characteristics
Now as we go to even higher chirality$(0.3)$, smaller regimes of self-propulsion expectedly give rise to mixed phases, as rotational tendencies of the particles dominate over their self-propulsion. On increasing self-propulsion we see particles are grouping to form small clusters with local cluster alignment similar to the case discussed previously. However, on continuing to increase self-propulsion there is a subtle difference from the previously discussed case of chirality with high self-propulsion. Now the cumulative effects of comparatively higher chirality and self-propulsion $v_{0}$ enhance the translational motion of the particles and deflect their trajectories, weakening the orientation interaction, in the process, resulting in a reduced density within the cluster (as shown in Fig.2II(l)). However, the chirality keeps the particles to be in a phase-separated state and maintains a common orientation with their neighbors as shown in Fig.3(l). As we continue to increase chirality to a value of $0.5$ particles exhibit mixed phases (as seen in Fig.2(I-II)(m) and Fig.3(m)) similar to the previously discussed cases for low self-propulsion velocities. Here on increasing self-propulsion, although the particles form local clusters, their enhanced intrinsic rotational dynamics destroy the overall ordering within the cluster in turn weakening their phase transition and ultimately leading to dilute and expanded clusters (as seen in Fig.2(I-II)(n-p) and Fig.3(n-p)). Based on the range of chirality $\Omega$ and activity $v_{0}$ we characterize the four cases in the system. We can define the ratio $r_{0}=\frac{v_{0}}{\Omega}$ as the radius of the trajectory of a single particle at the mean-field level. Hence the ratio determines the curvature of the particle trajectory. Large ratio $r_{0}$ means nearly straight line motion and small $r_{0}$ means a particle rotating about its own axis. The ratio $r_{0}$ can be compared to the one relevant length scale in the system: the interaction range $R$. Based on this comparison we can define the four scenarios in the parameter space.
@@
Case I: $r_{0}>>R$: In this case, particle trajectory is nearly a straight line within the range of its interaction radius, and two types of particles may not experience their chirality and for large $v_{0}$, they will remain mixed and try to move in the same direction (flocking) and global mixed ordered state will develop in the system. This is the situation for $(v_{0},\Omega)=(0.1-0.5,0.01)$ as can be seen in Fig. 2(I-II)(c-d) and 3(c-d). The cartoon of the trajectory of two types of particles is shown in Fig. 4(a).
@@
Case II: $r_{0}\geq R$: In this case, although particles maintain nearly the same orientation when lying within the interaction radius, very soon their trajectory gets deflected, they start moving in opposite directions and they leave the periphery of their previous neighbors and meet with other sets of neighbors. If the new neighbors have the same chirality, the particles will continue moving in the same direction. If not, their paths will be deflected again, leading them to a new location. This search will end with a state, where the majority of particles in the neighborhood are of similar chirality. Hence particles will phase separate, based on their chirality and macroscopic phase separation develops in the system. Each phase-separated cluster maintains ordering within the cluster as can be seen by Fig. 2(I-II)(g-h, l) and 3(g-h, l). The cartoon of the two particles’ trajectory and their interaction is shown in Fig. 4(b).
@@
Case III $r_{0}\simeq R$: In this case, the particle’s trajectory has larger curvature, and as soon as they leave the interaction radius they get separated from each other. But since the radius of trajectory is of the order of interaction radius, hence small pockets of clusters develop and microscopic clustering is obtained in the system. The mechanism of microscopic clustering is the same as explained in Case II with a key difference that now particles can only go to the distance smaller than that in Case II, this leads to the formation of small clusters, as can be seen by Fig. 2(I-II)(f, j) and 3(f, j). The cartoon of the trajectory of the two particles and their interaction is shown in Fig. 4(c).
@@
Case IV: $r_{0}<<R$: In this situation, particles do not get a chance to align and chiral nature dominates they keep rotating about their own axis and do not leave their interaction radius and hence remain mixed. as can be seen by Fig. 2(I-II)(i, m-n) and 3(i, m-n). The cartoon of the same is shown in Fig. 4(d).
@@
Till now we have discussed the characteristics of the system based on the snapshots, now we quantify the system behavior in the steady state by measuring the two order parameters, representing the phase separation and global orientation ordering. To explain the observed patterns as discussed before, we use the Phase separation order parameter $(PSOP)$ ($\Lambda$) to quantify the degree of mixing and demixing in the system and the Global order parameter $(GOP)$ ($\Phi$) to quantify the overall order and disorder within the system.
@@
where $\text{n}(\Omega^{+},t)$ and $\text{n}(\Omega^{-},t)$ represent the number of particles with positive and negative chirality, respectively at time $t$, within the distance $L/4$ (to differentiate the macroscopic phase separation in the system, We also checked for other distances, $L/16$, $L/8$, etc., the behaviour of $\Lambda$ remains the same.) Further, we take the mean of $\Lambda(t)$ over the steady state when the $\Lambda(t)$ remains statistically similar with respect to time. The mean is calculated over time in the steady state and over $10$ independent realizations.
@@
To further quantify the clustering and phase separation among the particles, we can also use the local density difference of two types of particles as defined before and shown in Fig. 2(II). Here we use the local density different $\Delta\rho({\bf r},t)$ at different times in the steady state. Using the local density different $\Delta\rho({\bf r},t)$ we calculated the two-point normalized correlation function $C(r,t)=<\Delta\rho({\bf r}_{0},t)\Delta\rho({\bf r}+{\bf r}_{0},t)>$, where mean is over all the points ${\bf r}_{0}$, directions, different times in the steady state and over $10$ ensembles. In the steady state the correlation function $C(r,t)$ will depend on the two parameters $v_{0}$ and $\Omega$. We further define the correlation length $\mathcal{L}(v_{0},\Omega,t)$ from the distance at which the correlation function reaches $10\%$ of its value from the first data.
@@
In Fig. 5(a-b) we show the plot of $PSOP$ and $\frac{\mathcal{L}(v_{0},\Omega)}{L/4}$ vs. $\Omega$ for different $v_{0}\in(0.01,0.5)$ respectively. Very clearly both show the non-monotonic trend as a function of chirality. For small activity $v_{0}=0.01$, particles have very weak self-propulsion speed and for any finite chirality, the rotational nature dominates, and both $PSOP$ and $\mathcal{L}(v_{0},\Omega)$ remain low with a small peak at $\Omega\simeq 0.02$. For very high $\Omega$, particles mainly remain stuck to their location and perform the circular motion. Hence do not get the chance to move out and phase separate. For moderate $\Omega$’s close to $0.02$, chirality and activity are in comparison and they have some tendency to form small clusters. On increasing activity, initially, the $PSOP$ and length $\mathcal{L}$ increases, and then again for larger chirality it shows the monotonic decay to small values. The $PSOP$ and $\mathcal{L}$ show the gradual increment on the increasing $v_{0}$ it reaches a maximum at $\Omega_{max}(v_{0})$. The $\Omega_{max}$ shifts towards the larger $\Omega$ on increasing $v_{0}$, although the shift is weak as shown in Fig. 5(b) (inset II). For activity ranging from medium to high ($v_{0}>0.05$), increasing the angular velocity $\Omega$ keeps the system mixed for very small $\Omega$, and both the $\Lambda$ and $\mathcal{L}$ are low, indicating a lack of phase separation. As the angular velocity $\Omega$ continues to increase, there is a bias in particle orientation, leading to an increase in phase separation. Consequently, the $\Lambda$ and $\mathcal{L}$ increase as the system and reaches to a maximum at $\Omega_{max}$ and maximum phase separation occurs with particles of different chiralities segregating into dense clusters, as depicted in Fig.2(I-II)(g-h, l). However, as $\Omega$ continue to increase, then beyond $\Omega_{max}$ particles do get separated, but their trajectory is small and they are separated in a sparser form, causing the $\Lambda$ and $\mathcal{L}$ to decrease gradually. Also, increasing activity $v_{0}$ leads to an increase in the phase separation as can be seen by the maximum value of $\Lambda_{max}$ increases with $v_{0}$. The plot of $\Lambda_{max}$ vs. $v_{0}$ is shown in Fig. 5(inset I). For lower values of $v_{0}$ $v_{0}\leq 0.05$, the system does not exhibit a phase-separated state, rather multiple small clusters are formed as can be seen by snapshots Fig. 2(I-II)(f, j) the small length $\Lambda$ and $\mathcal{L}$ values in Fig. 5(a-b). Consequently, the $\Lambda$ ($\mathcal{L}$) does not increase significantly as it does for higher $v_{0}$ values.
@@
Further the $\Phi(t)$ is averaged over the times in the steady state and ensembles to get the mean order parameter $\Phi$. The $\Phi$ is approximately equal to one when all particles move coherently in the same direction, indicating a highly ordered state. Conversely, the $\Phi$ approaches zero when the motion of individual particles is random, reflecting a disordered state.
@@
For activities in the medium to high range ($v_{0}\geq 0.1$), the global order parameter ($\Phi$) initially decreases from its maximum value as shown in Fig.5(c). At low values of $\Omega$, the $\Phi$ is at its peak because particles exhibit a low tendency to rotate and remain mixed and move in the same direction. Demixing creates the deflection in their trajectory and decreases in $\Phi$. The $\Phi$ also increases as $v_{0}$ increases while $\Omega$ remains low. This can be observed in the orientation snapshots of particles in Fig. 3(c-d), which shows that at $\Omega=0.01$, as $v_{0}$ increases, there is a higher likelihood that all particles within a cluster align in the same direction, resulting in a more uniform color within the cluster. This indicates that the particles have a reduced tendency to rotate, leading to more coherent movement in a single direction. Consequently, the system achieves a highly ordered state, which is reflected in Fig. 5(c), where the $\Phi$ reaches its maximum when $v_{0}$ is at its highest value under low $\Omega$ conditions.
@@
As $\Omega$ increases, the bias in particle orientation becomes stronger. This causes particles to rotate in different directions within a mixed state, resulting in the formation of distinct clusters containing both types of particles. In this mixed state, some clusters are dominated by particles with positive chirality, causing them to rotate clockwise $(CW)$, while others are dominated by particles with negative chirality, leading them to rotate anti-clockwise $(ACW)$ Fig. 5(c). However, these clusters are not perfectly ordered due to the mixed nature of the particles. As particles continuously join or leave the clusters, they do not rotate in a fixed radius, leading to frequent collisions between the mixed clusters. This behavior increases the overall randomness and disorder within the system, resulting in a sudden decrease in the $\Phi$ value, marking a significant change. This behavior is shown by particles when $v_{0}\geq 0.07$. At lower values of $v_{0}$, there is no sudden decrease in the $\Phi$; instead, it decreases slowly and becomes approximately constant at large $\Omega$, indicating that the particles are moving very slowly and there is no collision between particle clusters. The mechanism of collision is shown in the zoomed snapshot of the particles’ position with the orientation shown by arrows Fig. 5(c).
@@
However, the particles eventually become fully phase-separated at higher $v_{0}(\geq 0.1)$ and increasing $\Omega$. At high velocities, a second maxima of $\Phi$ is observed at $\Omega$ around $0.1$. Beyond this point, as $\Omega$ increases further, the $\Phi$ gradually decreases and stabilizes at an approximately constant value for a given $v_{0}$. Notably, the first maximum of the $\Lambda(\mathcal{L})$ and the second maximum of the $\Phi$ occur at the same value of $\Omega$ for a particular velocity, signifying the appearance of macroscopic phase separated state with global rotation of particles within their clusters in the system.
III.2 Dynamic of the system
Till now we discussed the characteristics of the system in terms of steady state in terms of ordering and density clusters for different activity and chirality. Now we will explore the dynamics of two types of particles in the system. We define the Mean Square displacement $(MSD)$ $\langle\Delta r^{2}\rangle(t)=\dfrac{1}{N}\langle\sum_{i=1}^{N}|\textbf{r}_{i}(t+t_{0})-\textbf{r}_{i}(t_{0})|^{2}\rangle$, where $\textbf{r}_{i}(t)$ is the position of the $i^{th}$ particle at time $t$, $t_{0}$ represents the reference time of the measurement. Here, .$\langle$....$\rangle$ … . . ⟩ indicates averaging over different values of $t_{0}$ in the steady state trajectory. In general, $MSD$ follows a power-law dependence on $t$. $\langle\Delta r^{2}\rangle\sim t^{\alpha}$, where the exponent $\alpha=1$ corresponds to diffusive motion, $\alpha<1$ to sub-diffusive motion, and $\alpha>1$ to super-diffusive motion, whereas for ballistic motion one has $\alpha=2$.

### v2

@@ III.1 Steady state characteristics
Now, as we go to even higher chirality $(0.3)$, smaller regimes of self-propulsion are expected to give rise to mixed phases as rotational tendencies of the particles dominate over their self-propulsion. On increasing self-propulsion, we see that the particles group to form small clusters with local cluster alignment, similar to the case discussed previously. However, there is a subtle difference from the previously discussed case of chirality with high self-propulsion in continuing to increase self-propulsion. Now, the cumulative effects of comparatively higher chirality and self-propulsion $v_{0}$ enhance the translational motion of the particles and deflect their trajectories, weakening the orientation interaction, in the process, resulting in a reduced density within the cluster (as shown in Fig.2II(l)). However, chirality keeps the particles in a phase-segregated state and maintains a common orientation with their neighbors, as shown in Fig.3(l). As we continue to increase chirality to a value of $0.5$, particles exhibit mixed phases (as seen in Figs.2(I-II)(m) and Fig.3(m)) similar to the previously discussed cases for low self-propulsion velocities. Here on increasing self-propulsion, although the particles form local clusters, their enhanced intrinsic rotational dynamics destroy the overall ordering within the cluster, in turn weakening their phase transition and ultimately leading to dilute and expanded clusters (as seen in Figs.2(I-II)(n-p) and Figs.3(n-p)).
Based on the range of chirality $\Omega$ and activity $v_{0}$, we characterize the four cases in the system. We can define the ratio $r_{0}=\frac{v_{0}}{\Omega}$ as the radius of the trajectory of a single particle at the mean-field level. Hence, the ratio determines the curvature of the particle trajectory. The large ratio $r_{0}$ means almost straight-line motion, and the small ratio $r_{0}$ means a particle rotating about its axis. The ratio $r_{0}$ can be compared to the one relevant length scale in the system: the interaction range $R$. Based on this comparison, we can define the four scenarios in the parameter space.
@@
Case I: $r_{0}>>R$: In this case, particle trajectory is nearly a straight line within the range of its interaction radius, and two types of particles may not experience their chirality, and for large $v_{0}$, they will remain mixed and try to move in the same direction (flocking), and a global mixed ordered state will develop in the system. This is the situation for $(v_{0},\Omega)=(0.1-0.5,0.01)$ as can be seen in Figs.2(I-II)(c-d) and 3(c-d). The cartoon of the trajectory of two types of particles is shown in Fig.4(a).
@@
Case II: $r_{0}\geq R$: In this case, although particles maintain nearly the same orientation when lying within the interaction radius, very soon their trajectory gets deflected, they start moving in opposite directions, and they leave the periphery of their previous neighbors and meet with other sets of neighbors. If the new neighbors have the same chirality, the particles will continue moving in the same direction. If not, their paths will be deflected again, leading them to a new location. This search will end with a state where most particles in the neighborhood are of similar chirality. Hence, particles will segregate phases based on their chirality, and a dense condensed phase develops in the system. Each phase segregates cluster maintains ordering within the cluster as can be seen by Figs.2(I-II)(g-h, l) and 3(g-h, l). The cartoon of the two particles’ trajectories and interactions is shown in Fig.4(b).
@@
Case III $r_{0}\lesssim R$: In this case, the particle’s trajectory has a larger curvature, and as soon as they leave the interaction radius, they segregate from each other. However, since the radius of the trajectory is of the order of the interaction radius, small pockets of clusters develop, and condensed clustering is obtained in the system. The mechanism of condensed clustering is the same as explained in Case II with a key difference that now particles can only go to a distance smaller than that in Case II; this leads to the formation of small clusters, as can be seen by Figs.2(I-II)(f,j) and 3(f,j). The cartoon of the trajectory of the two particles and their interaction is shown in Fig.4(c).
@@
Case IV: $r_{0}<<R$: In this situation, particles do not get a chance to align, and chiral nature dominates; they keep rotating about their axis and do not leave their interaction radius and hence remain mixed. As can be seen by Fig.2(I-II)(i, m-n) and 3(i, m-n). The cartoon of the same is shown in Fig.4(d).
Figure 5: (color online)$(a)$ Scaling of condensed polar packet density difference $(\Delta\rho^{*})$ with system size $(L)$ for four different parameters in the condensate phase. (bc)(b-c) show the saturation length $l_{sat}$ for the same micro and macro condensates respectively.
@@
We have discussed the system’s characteristics based on the snapshots; now, we quantify the system behavior in the steady state by measuring the two order parameters, representing the phase segregation and global orientation order. To explain the observed patterns, as discussed before, we use the Phase segregation order parameter $(PSOP)$ ($\Lambda$) to quantify the degree of mixing and demixing in the system, and the Global order parameter $(GOP)$ ($\Phi$) to quantify the overall order and disorder within the system.
@@
where $\text{n}(\Omega^{+},t)$ and $\text{n}(\Omega^{-},t)$ represent the number of particles with positive and negative chirality, respectively, at time $t$, within the distance $L/4$ (to differentiate the condensed phase segregation in the system. We also checked for other distances, $L/16$, $L/8$, etc., the behavior of $\Lambda$ remains the same.) Further, we take the mean of $\Lambda(t)$ over the steady state when the $\Lambda(t)$ remains statistically similar over time. The mean is calculated over time in the steady state and over $10$ independent realizations.
@@
To further quantify the clustering and phase segregation among the particles, we can also use the local density difference of the two types of particles as defined before and shown in Fig.2(II). Here, we use the local density difference $\Delta\rho({\bf r},t)$ at different times in the steady state. Using the local density difference $\Delta\rho({\bf r},t)$, we calculated the two-point normalized correlation function $C(r,t)=<\Delta\rho({\bf r}_{0},t)\Delta\rho({\bf r}+{\bf r}_{0},t)>$, where the mean is over all the points ${\bf r}_{0}$, directions, different times in the steady state, and over $10$ ensembles. The correlation function $C(r,t)$ will depend on the two parameters $v_{0}$ and $\Omega$ in the steady state. We further define the correlation length $\mathcal{L}(v_{0},\Omega,t)$ from the distance the correlation function reaches $10\%$ of its value from the first data point.
@@
In Figs.6(a-b) we show the plot of $PSOP$ and $\frac{\mathcal{L}(v_{0},\Omega)}{L/4}$ $vs.$ $\Omega$ for different $v_{0}\in(0.01,0.5)$ respectively. Very clearly, both show the non-monotonic trend as a function of chirality. For small activity $v_{0}=0.01$, particles have very weak self-propulsion speed, and for any finite chirality, the rotational nature dominates, and both $PSOP$ and $\mathcal{L}(v_{0},\Omega)$ remain low with a small peak at $\Omega\simeq 0.02$. For very high $\Omega$, particles mainly remain stuck to their location and perform a circular motion. Hence, they do not get the chance to move out and phase segregate. For moderate $\Omega$’s close to $0.02$, chirality and activity are in comparison, and they have some tendency to form small clusters. On increasing activity, initially, the $PSOP$ and length $\mathcal{L}$ increase, and then again, for larger chirality, it shows the monotonic decay to small values. The $PSOP$ and $\mathcal{L}$ show the gradual increment on the increasing $v_{0}$; it reaches a maximum at $\Omega_{max}(v_{0})$. The $\Omega_{max}$ shifts towards the larger $\Omega$ on increasing $v_{0}$, although the shift is weak, as shown in Fig.6(b) (inset II). For activity ranging from medium to high ($v_{0}>0.05$), increasing the angular velocity $\Omega$ keeps the system mixed for very small $\Omega$, and both the $\Lambda$ and $\mathcal{L}$ are low, indicating a lack of phase segregation. As the angular velocity $\Omega$ continues to increase, there is a bias in particle orientation, leading to an increase in phase segregation. Consequently, the $\Lambda$ and $\mathcal{L}$ increase as the system and reaches a maximum at $\Omega_{max}$ and maximum phase segregation occurs with particles of different chiralities segregating into dense clusters, as depicted in Figs.2(I-II)(g-h, l). However, as $\Omega$ continues to increase, the particles get segregated beyond $\Omega_{max}$. Still, their trajectory is small and segregated in a sparser form, causing the $\Lambda$ and $\mathcal{L}$ to decrease gradually. Also, increasing activity $v_{0}$ leads to an increase in the phase segregation, as can be seen by the maximum value of $\Lambda_{max}$ increasing with $v_{0}$. The plot of $\Lambda_{max}$ $vs.$ $v_{0}$ is shown in Fig.6(inset I). For lower values of $v_{0}$ $v_{0}\leq 0.05$, the system does not exhibit a phase segregated state, rather multiple small clusters are formed as can be seen by snapshots Figs.2(I-II)(f,j), the small length $\Lambda$ and $\mathcal{L}$ values in Figs.5(a-b). Consequently, the $\Lambda$ ($\mathcal{L}$) does not increase significantly as it does for higher $v_{0}$ values.
To further analyze the phase segregation on the variation of self-propulsion speed $v_{0}$. The effect of self-propulsion speed for different chirality is shown by the scaled plot as shown in Fig.6(d). We plot $\Lambda$ $vs.$ the scaled variable $v_{0}/\Omega$ for various fixed values of $\Omega$. For moderate to high chirality, all curves collapse onto a single master curve, revealing that the effective trajectory radius $r_{0}=v_{0}/\Omega$ is the key factor controlling phase segregation. A peak in $\Lambda$ occurs consistently near $v_{0}/\Omega\sim 1$, where the turning radius becomes comparable to the interaction range $R$, in agreement with the clustering mechanisms described in Cases II and III. For very small $\Omega$ (e.g. $0.01$ and $0.02$), this collapse fails, and $\Lambda$ remains low in all $v_{0}/\Omega$, indicating that particles behave nearly as straight movers and stay mixed. This confirms that $v_{0}/\Omega$ effectively captures the onset and strength of chirality-driven segregation in the system.
@@
Further, the $\Phi(t)$ is averaged over the times in the steady state and ensembles to get the mean order parameter $\Phi$. The $\Phi$ is approximately equal to one when all particles move coherently in the same direction, indicating a highly ordered state. Conversely, the $\Phi$ approaches zero when the motion of individual particles is random, reflecting a disordered state.
@@
For activities in the medium to high range ($v_{0}\geq 0.1$), the global order parameter ($\Phi$) initially decreases from its maximum value, as shown in Fig.6(c). At low values of $\Omega$, $\Phi$ is at its peak because the particles exhibit a low tendency to rotate and remain mixed and move in the same direction. Demixing creates a deflection in their trajectory and decreases in $\Phi$. The $\Phi$ also increases as $v_{0}$ increases while $\Omega$ remains low. This can be observed in the orientation snapshots of particles in Fig.3(c-d), which shows that at $\Omega=0.01$, as $v_{0}$ increases, there is a higher likelihood that all particles within a cluster align in the same direction, resulting in a more uniform color within the cluster. This indicates that the particles have a reduced tendency to rotate, leading to more coherent movement in a single direction. Consequently, the system achieves a highly ordered state, which is reflected in Fig.5(c), where the $\Phi$ reaches its maximum when $v_{0}$ is at its highest value under low $\Omega$ conditions.
@@
As $\Omega$ increases, the bias in particle orientation becomes stronger. This causes particles to rotate in different directions within a mixed state, forming distinct clusters that contain both types of particles. In this mixed state, some clusters are dominated by particles with positive chirality, causing them to rotate clockwise $(CW)$. In contrast, others are dominated by particles with negative chirality, leading them to rotate anticlockwise $(ACW)$ Fig.6(c). However, these clusters are not perfectly ordered due to the mixed nature of the particles. As particles continuously join or leave the clusters, they do not rotate at a fixed radius, leading to frequent collisions between the mixed clusters. This behavior increases the overall randomness and disorder within the system, resulting in a sudden decrease in the $\Phi$ value, marking a significant change. This behavior is shown by particles when $v_{0}\geq 0.07$. At lower values of $v_{0}$, there is no sudden decrease in the $\Phi$; instead, it decreases slowly and becomes approximately constant at large $\Omega$, indicating that the particles are moving very slowly and there is no collision between particle clusters. The mechanism of collision is shown in the zoomed snapshot of the particles’ position with the orientation indicated by arrows, Fig.5(c).
@@
However, particles eventually fully segregated phase at higher $v_{0}(\geq 0.1)$ and higher $\Omega$. At high velocities, a second maximum of $\Phi$ is observed at $\Omega$ around $0.1$. Beyond this point, as $\Omega$ further increases, $\Phi$ gradually decreases and stabilizes at an approximately constant value for a given $v_{0}$. In particular, the first maximum of $\Lambda(\mathcal{L})$ and the second maximum of the $\Phi$ occur at the exact value of $\Omega$ for a particular velocity, signifying the appearance of a dense condensed phase with global rotation of particles within their clusters in the system.
III.2 Dynamics of the system
We have now discussed the system’s characteristics regarding steady-state ordering and density clusters for different activities and chirality. Now, we will explore the dynamics of two types of particles in the system.